OpenStudy (hannahwaller5):

Will Give Medal!!! Two quadratic functions are shown. Function 1: f(x) = 3x2 + 6x + 7 Function 2: x g(x) −2 13 −1 7 0 3 1 7 Which function has the least minimum value and what are its coordinates? Function 1 has the least minimum value and its coordinates are (−1, 4). Function 1 has the least minimum value and its coordinates are (0, 7). Function 2 has the least minimum value and its coordinates are (−1, 7). Function 2 has the least minimum value and its coordinates are (0, 3).

OpenStudy (joshiscool):

Need Help? http://dotstorming.com/topic/570171aef45290eb0029bb89

OpenStudy (janu16):

when x=-2 f(-2)=3(-2)^2+6(-2)+7..=7 .and ....g(-2)=13 when x=-1 f(-1)=3(-1)^2+6(-1)+7....=4 and ...g(-1)=7 when x=0 f(0)=3(0)^2+6(0)+7..=7 ...and g(0)=3 when x=1 f(1)=3(1)^2+6(1)+7 ...= 16 and g(1)=7 so we compare to see which function has the least minimum value..

OpenStudy (janu16):

@Hannahwaller5

OpenStudy (hannahwaller5):

ok sorry

OpenStudy (janu16):

sow hcih function

OpenStudy (janu16):

which has the least minimum?

OpenStudy (janu16):

looks like function 2 but what coordinates

OpenStudy (janu16):

yes D

OpenStudy (hannahwaller5):

ok

OpenStudy (hannahwaller5):

Thanks lol im not good at math (: